# Direct Preference Optimization Using Sparse Feature-Level Constraints

### v1

@@ 4 Experimental Setup
Our model selection is guided by two key principles: scalability and transparency. For scalability, we first select a series of models spanning different parameter sizes, including Gemma-2-2B and Gemma-2-9B (Team et al., 2024) [^2] We select Gemma-scope as it provides pre-trained SAEs (Lieberum et al., 2024) for all layers.. This ensures that we can evaluate our approach’s performance as the model parameters scale and assess its robustness across diverse model architectures. For transparency, we exclusively select foundational models, which have not undergone supervised fine-tuning (SFT) or alignment processes. We begin by fine-tuning these models using a unified conversational format provided by the Halos dataset, applying it to the Ultrachat-200K (Ding et al., 2023). Dataset for initial instruction tuning. This establishes a baseline conversational capability and ensures that all our methods are compared on a consistent SFT model. Subsequently, we employ the UltraFeedback (Cui et al., 2024). Dataset to align the SFT models using various methods. This approach maintains transparency and control throughout the process, as all data and methods are open-sourced across the experimental setup.
@@
For analyzing the alignment and diversity trade-off of our method, following Zeng et al. (2024), in experiments, we validate and compare FPO against several strong alignment baselines, including DPO (Rafailov et al., 2024), SimPO (Meng et al., 2024), TDPO1, and TDPO2 Zeng et al. (2024).
@@ 5 Results and Discussions
Table 2: Performance comparison of different methods for Gemma-2-2B and Gemma-2-9B across various benchmarks (AlpacaEval-2, Arena-Hard, and MT-Bench), compared to Supervised Fine-Tuning (SFT), DPO and variants. Length controlled Winning Rate: WR-L; Winning Rate: WR.
Gemma-2-2B Gemma-2-9B Method AlpacaEval-2 Arena-Hard MT-Bench AlpacaEval-2 Arena-Hard MT-Bench FPO v.s. WR-L(%) WR (%) WR (%) $\Delta$ Score WR-L (%) WR (%) WR (%) $\Delta$ Score SFT 54.7 55.1 53.2 +0.5 51.2 52.4 53.4 +0.3 DPO 51.7 50.8 51.6 +0.1 51.0 51.0 51.2 +0.1 TDPO-1 51.5 54.4 51.4 +0.3 50.8 50.2 51.8 +0.1 TDPO-2 50.9 54.0 50.6 +0.2 50.2 49.9 49.5 0.0 SimPO 51.1 52.2 51.4 +0.4 50.2 51.8 51.0 +0.2
Table 3: Comparison of FPO and other baseline methods in terms of the trade-off between Alignment (accuracy) and Diversity (entropy) on the UltraFeedback dataset.
Method Accuracy (%) $\uparrow$ Diversity (Entropy) $\uparrow$
DPO 59.9 1.66
TDPO-1 63.2 1.65
TDPO-2 64.2 1.68
SimPO 63.4 1.64
FPO 64.1 1.68
5.1 The Trade-off between controllability and efficiency
@@
Search Strategy: Layer Selection layer $\ell$ 7 7 13 13 19 19 25 25 SAE type Residual MLP Residual MLP Residual MLP Residual MLP Accuracy (%) $\uparrow$ 57.2 57.4 59.1 61.3 59.7 62.4 63.6 63.4 Diversity (Entropy) $\uparrow$ 1.645 1.609 1.612 1.637 1.644 1.654 1.680 1.671 Search Strategy: $\alpha$ Selection / Stop-Gradient $\bm{\alpha}$ 0.1 0.5 1 2 0.1 0.5 1 2 Grad. sg. - - - - Chosen Chosen Chosen Chosen Accuracy (%) $\uparrow$ 64.1 63.7 63.4 61.9 64.0 63.6 62.7 62.1 Diversity (Entropy) $\uparrow$ 1.630 1.642 1.666 1.643 1.652 1.680 1.682 1.679
@@
Given the efficiency of FPO compared to TDPO2, as shown in the left one in Figure  4, we consider this result to be highly competitive. The efficiency of FPOis reflected primarily in two aspects: (1) Offline Processing. FPO does not require an additional reference model to be loaded during training, but only incurs minimal I/O overhead to read pre-stored information at each step, specifically the one-dimensional tensors needed for training. This process can be efficiently handled by the dataloader. (2) Sparsity. Due to the sparse activation values in the SAE encoder, we only need to process the activated values, reducing computational overhead. To validate its efficiency, we tested the memory consumption of different methods during training. In terms of memory usage, FPO maintains nearly the same level of memory consumption as reference-free methods like SimPO. Compared to methods that introduce more computation, such as TDPO, FPO achieves approximately a 17% memory optimization.
It is important to note that, compared to reference-free methods like SimPO, FPO still requires pre-computation of the reference model’s log probabilities and SAE feature activations. However, this reduces the peak computational and memory demands, making the model easier to run on smaller devices with lower costs. Considering that scaling up computational resources is generally more challenging than extending runtime, we believe this represents a reasonable trade-off between performance and cost. Additionally, it is worth noting that the hardware requirements during the caching process are significantly lower, as the model is only in inference mode.
Figure 3: Left. KL Divergence on the preferred responses (chosen). Center. KL Divergence on the dispreferred responses (rejected). Right. KL Divergence margin i.e., $|\beta D_{\text{SeqKL}}\left(x,y_{l};\pi_{\text{ref}}\|\pi_{\theta}\right)-\beta D_{\text{SeqKL}}\left(x,y_{w};\pi_{\text{ref}}\|\pi_{\theta}\right)|$.
Figure 4: Left. GPU memory consumption on a single H100 with all methods. We average the average GPU memory in 1,000 steps at the beginning of the training. Center. Feature-level MSE Loss of all methods after the whole alignment process. Here margin is defined as $|D^{\ell}_{\text{FPO}}\left(x,y_{l};\pi_{\text{ref}}\|\pi_{\theta}\right)-\beta D^{\ell}_{\text{FPO}}\left(x,y_{w};\pi_{\text{ref}}\|\pi_{\theta}\right)|$. Right. Win rates of FPO v.s. other methods above the improvements based on Gemma-2-2B evaluated by GPT-4 on different sampling temperatures.
@@ 5.2 Ablation Study
6 Related Work
Preference Optimization Methods in LLMs. Reinforcement learning (RL) has become a popular post-training technique, enabling models to learn implicit rewards from human feedback (Ouyang et al., 2022; Dubey et al., 2024; Yang et al., 2024a). The traditional RL pipeline involves training a reward model and updating the policy model via Proximal Policy Optimization (PPO) (Schulman et al., 2017). Recent work, such as DPO (Rafailov et al., 2024), leverages the log ratio between policy and target models to directly update policies based on the reward model’s objective. Extensions of DPO have introduced further refinements. KTO (Ethayarajh et al., 2024) eliminates pairwise data by modifying the value function using prospect theory, allowing training on individual sequences. Token-level DPO (Zeng et al., 2024) enforces constraints at the token level to improve generative diversity and also extends to the selection of specific tokens in pre-training (Lin et al., 2024) and post-training (Yang et al., 2024b). To reduce computational costs, ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024) remove the reference model, streamlining training. Our approach similarly omits the reference model for computational efficiency but uniquely integrates feature-level constraints to achieve both high efficiency and quality in preference learning.
Interpretating LLMs in Feature Space. One approach to LLM alignment focuses on enhancing transparency via mechanism interpretability (Shen et al., 2023; Wu et al., 2024). A central research goal in this area is understanding how LLMs internally extract, represent, and compute human-understandable features (Rai et al., 2024; Ferrando et al., 2024). Contrary to earlier assumptions, most neurons in LLMs do not activate exclusively for specific features but form polysemantic neurons (Mu & Andreas, 2020; Gurnee et al., 2023), a phenomenon termed ‘superposition’ (Elhage et al., 2022), which arises from compressing numerous learnable features into a limited number of dimensions (Hänni et al., 2024). Recent work shows that sparse autoencoders (SAE) can address this by decomposing internal representations into sparse, monosemantic features, improving interpretability (Huben et al., 2024; Templeton et al., 2024; Gao et al., 2024). Due to its scalability, SAE has been used to analyze LLM monosemanticity. Yan et al. (2024) found that alignment increases monosemanticity, while Marks et al. (2023) revealed that aligned LLMs learned feedback features related to human preferences, enhancing their output alignment. However, SAE has not yet been applied to construct feature-level constraints for improving LLM alignment.
@@ 7 Conclusion
In conclusion, we proposed FPO, a novel method for efficient and stable alignment of large language models using feature-level constraints. By leveraging sparse autoencoders and pre-computed offline references, FPO reduced the computational overhead traditionally associated with alignment methods like DPO and TDPO. Our experimental results demonstrate that FPO achieved significant improvements in alignment accuracy and diversity while maintaining low resource consumption. We prove that FPO achieved improvements over current state-of-the-art methods along all three dimensions: simplicity of implementation, efficiency, and generation quality.
Reproducible Statement
To ensure the reproducibility of our work, we have taken several steps to provide transparency in both the methodology and experimental setup. (1) All theoretical claims made in the paper are fully supported with proofs provided in the Appendix. We clearly state the assumptions and provide a detailed step-by-step explanation of our derivations to ensure clarity and completeness. (2) Details of the training process, including hyperparameters, optimizer settings, and batch sizes, are provided in Section 4 of the main text and the Appendix. We also describe the architecture of the models used, the pre-trained SAEs, and their configuration in detail. For hardware reproducibility, we have listed the GPUs used for each experiment and their respective memory consumption. (3) For reproducibility of the experiments, we utilize publicly available datasets. Detailed preprocessing steps for each dataset, including any data augmentation or filtering, are provided in the supplementary materials. Links to all datasets are included, and the steps to reproduce the exact test and training sets used in the paper are fully documented.(4) For our evaluation, we follow standard protocols and use well-established benchmarks such as AlpacaEval-2, Arena-Hard, and MT-Bench. The specific metrics and evaluation criteria are described in Section 5 of the paper, ensuring consistency and repeatability across different model sizes and configurations.
By ensuring all aspects of the work are thoroughly documented and available, we strive to make the replication of our results straightforward and accessible to the research community.
@@ Appendix A Training Settings
Table 5: Hyperparameters for Gemma-2-2b and Gemma-2-9b
@@ Appendix B Bounding KL Divergence with MSE of Sparse Activation
Definition 1 (Sparse Activations).
@@
Definition 2 (Token Logits and Probabilities).

### v4

@@ 4 Experimental Setup
Our model selection is guided by two key principles: scalability and transparency. For scalability, we first select a series of models spanning different parameter sizes, including Gemma-2-2B and Gemma-2-9B (Team et al., 2024) [^2] We select Gemma-scope as it provides pre-trained SAEs (Lieberum et al., 2024) for all layers.. This ensures that we can evaluate our approach’s performance as the model parameters scale and assess its robustness across diverse model architectures.
For transparency, we exclusively select foundational models, which have not undergone supervised fine-tuning (SFT) or alignment processes. We begin by fine-tuning these models using a unified conversational format provided by the Halos dataset, applying it to the Ultrachat-200K (Ding et al., 2023). Dataset for initial instruction tuning. This establishes a baseline conversational capability and ensures that all our methods are compared on a consistent SFT model. Subsequently, we employ the UltraFeedback (Cui et al., 2024). Dataset to align the SFT models using various methods. This approach maintains transparency and control throughout the process, as all data and methods are open-sourced across the experimental setup.
@@
Figure 3: Left 1. KL Divergence on the preferred responses (chosen). Left 2. KL Divergence on the dispreferred responses (rejected). Right 1. KL Divergence margin i.e., $|\beta D_{\text{SeqKL}}\left(x,y_{l};\pi_{\text{ref}}\|\pi_{\theta}\right)-\beta D_{\text{SeqKL}}\left(x,y_{w};\pi_{\text{ref}}\|\pi_{\theta}\right)|$. Right 2. Win rates of FPO v.s. other methods above the improvements based on Gemma-2-2B on different sampling temperatures.
@@
For analyzing the alignment and diversity trade-off of our method, following Zeng et al. (2024), in experiments, we validate and compare FPO against several strong alignment baselines, including DPO (Rafailov et al., 2024), SimPO (Meng et al., 2024), TDPO1, and TDPO2 (Zeng et al., 2024).
@@ 5 Results and Discussions
5.1 The Trade-off between Controllability and Efficiency.
@@
Search Strategy: Layer Selection Search Strategy: $\alpha$ Selection / Stop-Gradient
layer $\ell$ 7 7 13 13 19 19 25 25 $\boldsymbol{\alpha}$ 0.1 0.5 1 2 0.1 0.5 1 2
SAE type Res MLP Res MLP Res MLP Res MLP Grad sg. - - - - Yes Yes Yes Yes
Acc (%) $\uparrow$ 57.2 57.4 59.1 61.3 59.7 62.4 63.6 63.4 64.1 63.7 63.4 61.9 64.0 63.6 62.7 62.1
$H\uparrow$ 1.645 1.609 1.612 1.637 1.644 1.654 1.680 1.671 1.630 1.642 1.666 1.643 1.652 1.680 1.682 1.679
@@
Given the efficiency of FPO compared to TDPO2, as shown in the left one in Figure  4, we consider this result to be highly competitive. The efficiency of FPOis reflected primarily in two aspects: (1) Offline Processing. FPO does not require an additional reference model to be loaded during training, but only incurs minimal I/O overhead to read pre-stored information at each step, specifically the one-dimensional tensors needed for training. This process can be efficiently handled by the dataloader. (2) Sparsity. Due to the sparse activation, we only need to process the activated values, reducing computational overhead. To validate its efficiency, we tested the memory consumption of different methods during training. In terms of memory usage, FPO maintains nearly the same level of memory consumption as reference-free methods like SimPO. Compared to methods that introduce more computation, such as TDPO, FPO achieves approximately a 17% memory optimization.
It is important to note that, compared to reference-free methods like SimPO, FPO still requires pre-computation of the reference model’s log probabilities and SAE feature activations. However, this reduces the peak computational and memory demands, making the model easier to run on smaller devices with lower costs. Considering that scaling up computational resources is generally more challenging than extending runtime, we believe this represents a reasonable trade-off between performance and cost.
Figure 4: Left. GPU memory consumption on a single H100 with all methods. We average the average GPU memory in 1,000 steps at the beginning of the training. Right. Feature-level MSE Loss of all methods after the whole alignment process. Here margin is defined as $|D^{\ell}_{\text{FPO}}\left(x,y_{l};\pi_{\text{ref}}\|\pi_{\theta}\right)-\beta D^{\ell}_{\text{FPO}}\left(x,y_{w};\pi_{\text{ref}}\|\pi_{\theta}\right)|$. The close correspondence between the MSE Loss margin reduction and KL divergence margin reduction supports the validity of our approach.
@@ 5.2 Ablation Study
5.3 FPO Achieves Accurate Control Over Model Capabilities
A key advantage of FPO lies in its ability to precisely control model capabilities. While efficiency and reduced memory usage are significant outcomes, the accurately in controlling model behavior stems from FPO’s underlying mechanism.
Table 4: FPO Accurate Control Experiment Highlights
Domain Capability Method/Setting Metric Value
Instruction Following JSON Format FPO ($\beta=0$ on all format features) Accuracy 0.46 / 0.00
FPO ($\beta=1$ on all format features) 0.03 / 0.00
FPO ($\beta=1$ on JSON feature) 0.04 / 0.00
Multilingual French FPO ($\beta=0$ on French feature) Output Rate % 80
FPO ($\beta=1$ on French feature) 3
Safety - FPO ($\beta=0$) Attack Success Rate % 30
FPO ($\beta=1$ on safety-related features) 80
FPO ($\beta=1$ on harmful features) 5
Sentiment Positive FPO ($\beta=0$) Positive Sentiment Ratio 65
FPO ($\beta=1$ on Positive feature) 5
To substantiate this, experiments were conducted to accurately regulate specific capabilities during alignment, while leaving others unaffected and maintaining strong overall performance. These experiments assessed four critical domains: (1) Instruction Following: Evaluated using IFEval, focusing on tasks such as JSON formatting, capitalizing, highlighting text, lowercasing, creating bullet lists, and using quotations. (2) Multilingual Capability: Assessed with MultiAlpaca and WildChat datasets, covering French, Canadian French, German, and Italian, with English questions from MKQA. (3) Safety: Measured using Jailbreak Bench and AdvBench. (4) Sentiment: Analyzed with the Twitter Financial News Sentiment dataset.
The experimental setup involved the Gemma2-2B model with an SAE width of 16k. Features relevant to each domain were identified by first selecting the top 50 features most aligned with each target domain based on single-token activations, followed by validating their global relevance through average activations across target datasets. The model was trained with a learning rate of 2e-5, a batch size of 64, and a maximum sequence length of 2048, using an Adam optimizer and a cosine learning rate schedule with 10% warmup steps over 1 epoch. Hyperparameters such as $\alpha$ for TDPO2 and FPO were set to 0.5, $\beta$ for SimPO to 2, and $\beta$ for other methods to 0.1.
The results demonstrate FPO’s capacity for targeted control. For example, in Instruction Following, by adjusting the $\beta$ value for all format-related features to 1, FPO significantly reduced the model’s propensity to use these formats (e.g., JSON format accuracy dropped to 0.03 with instruction and 0.00 without, compared to 0.46 and 0.00 respectively when $\beta=0$). Conversely, setting $\beta=0$ for these features (effectively removing the constraint) maintained or enhanced these abilities (e.g., JSON format accuracy of 0.46 with instruction). When targeting a specific feature like JSON formatting by setting its $\beta=1$, the JSON format accuracy dropped to 0.04 (with instruction), while other formatting abilities like capitalizing (0.73) or highlighting (0.55) remained high.
Similar precise control was observed in other domains. For Multilingual Capability, setting $\beta=1$ for French-related features drastically reduced French output (3% output rate) while other languages like German (45%) and Italian (50%) were less affected compared to when $\beta=0$ for French features (French output rate of 80%). In Safety, applying $\beta=1$ to safety-related features increased the attack success rate to 80% (indicating reduced safety), whereas applying it to harmful features decreased the attack success rate to 5% (indicating enhanced safety), compared to a 30% success rate when $\beta=0$. For Sentiment, setting $\beta=1$ on positive sentiment features reduced positive sentiment expression to 5% and increased negative sentiment to 95%. Conversely, targeting negative sentiment features with $\beta=1$ resulted in 93% positive and 7% negative sentiment outputs.
@@ 6 Conclusion
In conclusion, we proposed FPO, a novel method for efficient and stable alignment of large language models using feature-level constraints. By leveraging sparse autoencoders and pre-computed offline references, FPO reduced the computational overhead traditionally associated with alignment methods like DPO and TDPO. Our experimental results demonstrate that FPO achieved significant improvements in alignment accuracy and diversity while maintaining low resource consumption.
Impact Statement
This paper proposes Feature-level Constrained Direct Preference Optimization (FPO) to address the challenges of computational inefficiency and training instability in aligning large language models (LLMs) with human preferences. Its impact spans multiple aspects. In academic research, FPO offers a novel perspective by being the first to integrate sparse feature-level constraints into LLM alignment, inspiring further exploration in this area. Methodologically, it enriches the field by using Sparse Autoencoders (SAEs) to approximate KL divergence, balancing efficiency and controllability. On the industry front, FPO promotes the safe and reliable application of LLMs in critical sectors, such as healthcare, finance, and education. Its use of open-source models and datasets fosters the development of the open-source ecosystem, encouraging collaboration and innovation among researchers.
Acknowledgement
This publication has been supported by the National Natural Science Foundation of China (NSFC) Key Project under Grant Number 62336006.
@@ Appendix A Training Settings
Table 5: Hyperparameters for Gemma-2-2b and Gemma-2-9b.
@@ Appendix B Bounding KL Divergence with MSE of Sparse Activation
Definition B.2 (Sparse Activations).
@@
Definition B.3 (Token Logits and Probabilities).
@@
Appendix C Concrete Examples of Feature-Level Representations vs. Token-Level Embeddings
This section provides concrete examples and visualizations to highlight the differences between feature-level representations and token-level embeddings in our framework.
C.1 Definitions and Intuitions
Token-Level Embeddings:
Token-level embeddings correspond directly to the token output probabilities (logits) generated by a model. These embeddings are high-dimensional vectors representing each token in the model’s vocabulary. For a sequence $x=[x_{1},x_{2},\dots,x_{T}]$, the token-level embeddings at position $t$ are computed as:
$h_{t}=f_{\text{token}}(x_{t})\in\mathbb{R}^{V},$
where $V$ is the vocabulary size, and $f_{\text{token}}$ is the output projection from the model’s hidden state.
Feature-Level Representations:
Feature-level representations, on the other hand, are high-level abstractions derived from the model’s intermediate layers. These representations capture patterns and salient features across sequences. Using a Sparse Autoencoder (SAE), the hidden state $h_{t}^{\ell}$ at layer $\ell$ can be transformed into sparse activations $c_{t}^{\ell}$, defined as:
$c_{t}^{\ell}=\text{ReLU}(W_{\text{enc}}h_{t}^{\ell}+b),$
where $W_{\text{enc}}\in\mathbb{R}^{m\times d}$, $b\in\mathbb{R}^{m}$, and $m\ll V$. This sparse activation ensures only a subset of features is active, making the representation interpretable and efficient.
C.2 Concrete Example: A Mathematical Query
Consider the input query:
”What is the derivative of $x^{2}+3x+5$?”
Token-Level Embedding:
The token-level output probabilities for each token in the response sequence, such as ”The derivative is 2x + 3.”, involve logits for every token:
[…]
Feature-Level Representation:
Using SAE on the 25th layer, the sparse feature representation for the same sequence might activate specific features corresponding to mathematical operations or semantic groupings:
$c^{\ell}=[\text{activation}_{1}(\text{Polynomial}),\text{activation}_{2}(\text{Arithmetic}),\dots].$
Appendix D Experiments on Additional Baselines and Ablation Studies
In response to reviewer feedback, we conducted additional experiments to address their concerns and validate our methodology. These include comparisons with the SimPO+KL baseline and ablations on multi-layer sparse autoencoders (SAEs).
D.1 Comparison with SimPO+KL
This subsection provides a direct comparison of our method against SimPO+KL. We implemented SimPO+KL following the same experimental settings in Section 4. Specifically, we tested on the Gemma-2-2B model using the AlpacaEval-2 dataset, evaluating both winning rate (WR) and length-controlled winning rate (WR-L). Results are summarized in Table 6.
Table 6: Comparison of FPO with SimPO+KL on the AlpacaEval-2 dataset. Metrics include Accuracy (%), Diversity (Entropy), WR (%), and WR-L (%).
Method Accuracy (%) $\uparrow$ Diversity (Entropy) $\uparrow$ WR (%) $\uparrow$ WR-L (%) $\uparrow$
FPO (Ours) 64.1 1.68 51.8 50.2
SimPO+KL 63.6 1.66 50.8 50.6
SimPO 63.4 1.64 50.2 49.8
TDPO-2 64.2 1.68 50.0 50.0
Discussion:
The results show that FPO achieves comparable or better performance than SimPO+KL in both WR and WR-L metrics. This highlights the effectiveness of feature-level constraints in maintaining both alignment quality and diversity, with a competitive computational cost.
D.2 Ablation Study on Multi-Layer SAEs
To find out the effect of extending SAEs across multiple layers, we conducted experiments adding SAEs at different layer combinations. Table 7 presents the performance metrics when SAEs were applied to various combinations of shallow, middle, and deep layers.
Table 7: Ablation study on using SAEs at multiple layers in FPO. Metrics include Accuracy (%), Diversity (Entropy), WR (%), and WR-L (%).
SAE Layers Accuracy (%) $\uparrow$ Diversity (Entropy) $\uparrow$ WR (%) $\uparrow$ WR-L (%) $\uparrow$
Single Layer (Layer 25) 64.1 1.68 51.8 50.2
Layers 0, 25 62.1 1.70 47.2 48.8
Layers 12, 25 61.9 1.64 48.4 49.5
Layers 24, 25 58.2 1.66 48.6 46.4
Layers 0, 12, 25 51.4 1.66 47.8 49.8
Discussion:
Results indicate that adding multiple SAE layers does not consistently improve performance and may even degrade alignment metrics (e.g., accuracy and WR). The best results were achieved with a single SAE layer (Layer 25), confirming that simplicity in feature extraction leads to more stable alignment.